# Dual process in the two-parameter Poisson–Dirichlet diffusion

### v5

@@ 5 Derivation of the transition density
In this section we apply the results of the previous two sections, namely the identified dual process together with conditional two-parameter partition structures, to derive the transition density of the two-parameter diffusion. In preparation to this task, recall that the two-parameter diffusion is reversible with stationary distribution $\mathrm{PD}_{\alpha,\theta}$. The transition probability $P(t,x,dy)$ is absolutely continuous with respect to $\mathrm{PD}_{\alpha,\theta}$, and its transition density $p(t,x,y)$ was shown in [29] to be
@@
The proof of (24) in [29] leverages on a spectral representation by expanding on a proof by [17] for the infinitely-many-alleles model (the one-parameter case). Previously, [39] had obtained (24) in the one-parameter case as a limit from a model with finitely-many types (cf. also Proposition 4.3 in [44]), and [19] the corresponding version for the labelled model. Analogous transition structures also appear in a family of diffusions defined through the Jack graph, see [72] and references therein.
Here it is useful to point out that there are two possible expansions for the transition density of the model at hand: a spectral expansion in terms of reproducing kernel orthogonal polynomials on $\mathrm{PD}_{\alpha,\theta}$, and an expansion as mixture of two-parameter Poisson–Dirichlet distributions. The equivalence of the two forms was explained in [43, 45] for the one parameter case and in [71] (cf. Theorem 2.1) for the two-parameter case. In particular, [71] showed that (24) is the same as
@@
where ${\tilde{d}_{1}^{\theta}}(t)=1-\sum_{n=2}^{\infty}d_{n}^{\theta}(t)=d_{0}^{\theta}(t)+d_{1}^{\theta}(t)>0$. Note that algebraically ${\tilde{d}_{1}^{\theta}}(t)=d_{0}^{\theta}(t)+d_{1}^{\theta}(t)$ with both $d_{0}^{\theta}(t),d_{1}^{\theta}(t)$ non-negative when $\theta\geq 0$, but for $-1<\theta<0$ these may not be individually non-negative.
@@
(26) $\mathbb{E}_{x}\big{[}{P}_{\eta}\big{(}X(t)\big{)}\big{]}=\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}\Bigg{(}\tilde{d}^{\theta}_{n1}(t)+\sum_{w=2}^{n}d^{\theta}_{nw}(t)\sum_{|\omega|=w,\omega\subset\eta}{\cal H}(\omega|\eta)\frac{{P}_{\omega}(x)}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}\Bigg{)},$
@@
The proof is immediate by expanding the left hand side of (13) and using (19). ∎
Note that in light of (5), we can replace ${P}_{\eta}$ and ${P}_{\omega}$ with $\vec{P}_{\eta}$ and $\vec{P}_{\omega}$ in (26). In this equation there are three types of sampling probabilities, namely
@@
Here $M_{n}(\eta)$ is as in (8) and it is interpreted as the law of the partition $\eta$ taken at stationarity; $M_{n,t}(\eta)$ is the law of $\eta$ relative to the conditional distribution of $X(t)$, given $X(0)=x$; finally $M_{n,\omega}(\eta)$ is the law of $\eta$ conditional of having observed $\omega\subset\eta$, which was the object of Section 4 (cf. (23)). Since all three are partition structures, we can define the measures
@@
Recall from Section 4 that $\pi(X^{n-w},Y^{w})$ denotes the partition of $n$ induced by the joint sample $(X^{n-w},Y^{w})$. By Kingman’s representation theorem (cf. Section 2), $\pi(X^{n-w},Y^{w})\big{/}n$ converges almost surely to a random variable $Z$ with values in $\overline{\nabla}_{\infty}$. Because the conditional distribution of $\pi(X^{n-w},Y^{w})\big{/}n$ is
@@
we know that $\nu_{n,\omega}(dy)$ converges weakly to the conditional distribution $\mathbb{P}(Z\in dy|Y^{w})$. Conditioning on $Y^{w}$, we know $\pi(X^{n-w})\big{/}n$ and $\pi(X^{n-w},Y^{w})\big{/}n$ will both converge to $Z$ almost surely. Then $Z$ has distribution $\mathrm{PD}_{\alpha,\theta|\omega}$ due to Lemma 2.1 and equation (22). ∎
@@
Given $X(0)=x$, $X(t)$ has a distribution given by the mixture
(28) $P(t,x,dy)=\widetilde{d}_{1}^{\theta}(t)\mathrm{PD}_{\alpha,\theta}(dy)+\sum_{w=2}^{\infty}d^{\theta}_{w}(t)\sum_{\omega:|\omega|=w}\mathrm{PD}_{\alpha,\theta|\omega}(dy)\vec{P}_{\omega}(x).$
@@
From Lemma 5.1 we have
$\mu_{n,t,x}(dy)=$ $\tilde{d}_{n,1}^{\theta}(t)\mu_{n}(dy)+\sum_{w=2}^{n}d^{\theta}_{nw}(t)\sum_{|\omega|=w}\vec{P}_{\omega}(x)\nu_{\omega,n}(dy).$
@@
¯$\int_{\overline{\nabla}_{\infty}}f(y)\mu_{n,t,x}(dy)=\tilde{d}_{n,1}^{\theta}(t)\int_{\overline{\nabla}_{\infty}}f(y)\mu_{n}(dy)+\sum_{w=2}^{n}d^{\theta}_{nw}(t)\sum_{|\omega|=w}\vec{P}_{\omega}(x)\int_{\overline{\nabla}_{\infty}}f(y)\nu_{\omega,n}(dy).$
@@
$\,+\sum_{w=2}^{\infty}d^{\theta}_{w}(t)\sum_{|\omega|=w}\vec{P}_{\omega}(x)\int_{\overline{\nabla}_{\infty}}f(y)\mathrm{PD}_{\alpha,\theta|\omega}(dy).$
@@
$P(t,x,dy)=\tilde{d}_{1}^{\theta}(t)\mathrm{PD}_{\alpha,\theta}(dy)+\sum_{w=2}^{\infty}d^{\theta}_{w}(t)\sum_{|\omega|=w}\vec{P}_{\omega}(x)\mathrm{PD}_{\alpha,\theta|\omega}(dy)$
@@
The transition function obtained in Proposition 5.3 can be interpreted in terms of Pólya urn schemes. More specifically, we can exploit the the Pólya urn representation discussed in Section 4 to construct, for each $t\geq 0$, a random variable $X(t)$ with distribution $P(t,x,dy)$ as in Proposition 5.3. In short, this is obtained by branching the urn scheme after the observation of a sequence that generates a configuration $\omega$, and letting the two resulting split urns evolve independently onwards. This approach is inspired by a similar construction in [44] for the one-parameter case (see also [42] for a connection with Wright–Fisher diffusion bridges).
Let $\{U_{n},n\geq 1\}$ be the process driven by the Pólya urn scheme described in Section 4, where the urn starts with a single white ball and $U_{n}$ is the configuration of balls after $n$ draws. The urn scheme is run for a random number of draws $D_{t}$ to obtain a configuration $U_{D_{t}}$, where $\mathbb{P}(D_{t}=w)=d_{w}^{\theta}(t),~{}w\geq 1,$ and $D_{t}$ is as in (15) and $d_{w}^{\theta}(t)$ as in (17). The urn is then split and two urns are run independently for $n-D_{t}$ additional draws, both beginning at $U_{D_{t}}$. This produces two random partitions $\eta$ and $\tilde{\eta}$, induced by the two resulting urn configurations. Conditional on the partition $\omega$ obtained at step $D_{t}$, these partitions are independent and with common distribution (23), namely
@@
$\mathbb{P}(U_{n}=\eta,U^{\prime}_{n}=\eta^{\prime})=\sum_{w=1}^{\infty}d_{w}^{\theta}(t)\sum_{|\omega|=w}\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}\mathbb{P}(\eta|\omega)\mathbb{P}(\eta^{\prime}|\omega),$
@@
$\left(\sum_{w=1}^{\infty}d_{w}^{\theta}(t)p_{w}(x,y)\right)\mathrm{PD}_{\alpha,\theta}(dx)\mathrm{PD}_{\alpha,\theta}(dy).$
This probabilistic construction shows that the conditional distribution of $Y$ given $X$ is the same as the transition density (25) of the diffusion, at a fixed time $t$.
@@ Appendix A Appendix
by simply applying $L_{\alpha,\theta}$. If $a_{1}(\eta)>0$ $P_{\eta}$ can be recursively expressed as linear combinations of functions in ${\cal C}$ which can then be acted on by the differential form (1). Recall that $P_{\eta}$ is exchangeable in the elements of $\eta$ so it is possible to rearrange the elements so that the $a_{1}(\eta)$ singletons are the last entries of $\eta$. Suppose that $\eta_{d}=1$. Then we use the notation that $\eta^{-}=\eta-e_{d}=(\eta_{1},\ldots,\eta_{d-1})$, where the final component of $\eta$ is removed. Consider (9) with $i=d$, then
@@
MR acknowledges support of MUR - Prin 2022 - Grant no. 2022CLTYP4, funded by the European Union - Next Generation EU. YZ was supported by grant NSFC117015170.

### v6

@@ 5 Derivation of the transition density
In this section we apply the results of the previous two sections, namely the identified dual process together with conditional two-parameter partition structures, to derive the transition density of the two-parameter diffusion. In preparation to this task, recall that the two-parameter diffusion is reversible with stationary distribution $\mathrm{PD}_{\alpha,\theta}$ [60]. The transition probability $P(t,x,dy)$ is absolutely continuous with respect to $\mathrm{PD}_{\alpha,\theta}$, and its transition density $p(t,x,y)$ was shown in [29] to be
@@
The proof of this fact in [29] leverages on a spectral representation by expanding on a proof by [17] for the the one-parameter case, i.e., the infinitely-many-neutral-alleles model. Previously, [38] had obtained (25) in the one-parameter case as a limit from a model with finitely-many types (cf. also Proposition 4.3 in [43]), and [19] the corresponding version for the labelled model. Analogous transition structures also appear in a family of diffusions defined through the Jack graph, see [70] and references therein.
Here it is useful to emphasize that there are two possible expansions for the transition density of the model at hand: a spectral expansion in terms of reproducing kernel orthogonal polynomials on $\mathrm{PD}_{\alpha,\theta}$, and an expansion as mixture of two-parameter Poisson–Dirichlet distributions. The equivalence of the two forms was explained in [42, 44] for the one parameter case and in [69] (cf. Theorem 2.1) for the two-parameter case. In particular, [69] showed that (25) is the same as
@@
where
(27) $\tilde{d}_{1}^{\theta}(t):=1-\sum_{n=2}^{\infty}d_{n}^{\theta}(t).$
Note that algebraically $\tilde{d}_{1}^{\theta}(t)=d_{0}^{\theta}(t)+d_{1}^{\theta}(t)$ with both $d_{0}^{\theta}(t),d_{1}^{\theta}(t)$ non-negative when $\theta\geq 0$, but for $-1<\theta<0$ these may not be individually non-negative.
@@
(28) $\mathbb{E}_{x}\big{[}{P}_{\eta}\big{(}X(t)\big{)}\big{]}=\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\eta}\big{]}\Bigg{(}\tilde{d}_{n1}^{\theta}(t)+\sum_{\ell=2}^{n}d^{\theta}_{n\ell}(t)\sum_{|\omega|=\ell,\omega\subset\eta}{\cal H}(\omega|\eta)\frac{{P}_{\omega}(x)}{\mathbb{E}_{\alpha,\theta}\big{[}{P}_{\omega}\big{]}}\Bigg{)},$
with $\tilde{d}_{n1}^{\theta}(t)$ defined as in (27) in relation to (17).
@@
The proof is immediate by expanding the right-hand side of (14) and using (20). ∎
Note that in light of (5), we can replace ${P}_{\eta}$ and ${P}_{\omega}$ with $\vec{P}_{\eta}$ and $\vec{P}_{\omega}$ in (28). In this equation there are three sampling probabilities, namely
@@
Here $M_{n}(\eta)$ is as in (8) and it is interpreted as the law of the partition $\eta$ taken at stationarity; $M_{n,t}(\eta)$ is the law of $\eta$ relative to the conditional distribution of $X(t)$, given the initial state $X(0)=x$; finally $M_{n,\omega}(\eta)$ is the law of $\eta$ conditional of having observed $\omega\subset\eta$, which was the object of Section 4 (cf. (24)). Since all three are partition structures, we can define the respective associated measures
@@
Recall from Section 4 that $\pi(X^{n-\ell},Y^{\ell})$ denotes the partition of $n$ induced by the joint sample $(X^{n-\ell},Y^{\ell})$. By Kingman’s representation theorem (cf. Section 2), $\pi(X^{n-\ell},Y^{\ell})\big{/}n$ converges almost surely to a random variable $Z$ with values in $\overline{\nabla}_{\infty}$. Because the conditional distribution of $\pi(X^{n-\ell},Y^{\ell})\big{/}n$ is
@@
we know that $\nu_{n,\omega}(dy)$ converges weakly to the conditional distribution $\mathbb{P}(Z\in dy|Y^{\ell})$. Conditioning on $Y^{\ell}$, we know $\pi(X^{n-\ell})\big{/}n$ and $\pi(X^{n-\ell},Y^{\ell})\big{/}n$ will both converge to $Z$ almost surely. Then $Z$ has distribution $\mathrm{PD}_{\alpha,\theta|\omega}$ due to Lemma 2.1 and equation (23). ∎
@@
Given $X(0)=x$, $X(t)$ has distribution
(30) $P(t,x,dy)=\tilde{d}_{1}^{\theta}(t)\mathrm{PD}_{\alpha,\theta}(dy)+\sum_{\ell=2}^{\infty}d^{\theta}_{\ell}(t)\sum_{\omega:|\omega|=\ell}\mathrm{PD}_{\alpha,\theta|\omega}(dy)\vec{P}_{\omega}(x),$
with $\tilde{d}_{1}^{\theta}(t)$ as in (27).
@@
From Corollary 5.1 we have
$\mu_{n,t,x}(dy)=$ $\tilde{d}_{n,1}^{\theta}(t)\mu_{n}(dy)+\sum_{\ell=2}^{n}d^{\theta}_{n\ell}(t)\sum_{|\omega|=\ell}\vec{P}_{\omega}(x)\nu_{\omega,n}(dy).$
@@
¯$\int_{\overline{\nabla}_{\infty}}f(y)\mu_{n,t,x}(dy)=\tilde{d}_{n,1}^{\theta}(t)\int_{\overline{\nabla}_{\infty}}f(y)\mu_{n}(dy)+\sum_{\ell=2}^{n}d^{\theta}_{n\ell}(t)\sum_{|\omega|=\ell}\vec{P}_{\omega}(x)\int_{\overline{\nabla}_{\infty}}f(y)\nu_{\omega,n}(dy).$
@@
$\,+\sum_{\ell=2}^{\infty}d^{\theta}_{\ell}(t)\sum_{|\omega|=\ell}\vec{P}_{\omega}(x)\int_{\overline{\nabla}_{\infty}}f(y)\mathrm{PD}_{\alpha,\theta|\omega}(dy).$
@@
$P(t,x,dy)=\tilde{d}_{1}^{\theta}(t)\mathrm{PD}_{\alpha,\theta}(dy)+\sum_{\ell=2}^{\infty}d^{\theta}_{\ell}(t)\sum_{|\omega|=\ell}\vec{P}_{\omega}(x)\mathrm{PD}_{\alpha,\theta|\omega}(dy)$
@@
The transition function obtained in Proposition 5.3 can be interpreted in terms of Pólya urn schemes. More specifically, we can exploit the Pólya urn representation discussed in Section 4 to construct, for each $t\geq 0$, a random variable $X(t)$ with distribution $P(t,x,dy)$ as in Proposition 5.3. In short, this is obtained by branching the urn scheme after the observation of a sequence that generates a configuration $\omega$, and letting the two resulting split urns evolve independently onwards. This approach is inspired by a similar construction in [43] for the one-parameter case (see also [41] for a connection with Wright–Fisher diffusion bridges).
Let $\{U_{n},n\geq 1\}$ be the process driven by the Pólya urn scheme described in Section 4, where the urn starts with a single white ball and $U_{n}$ is the configuration of balls after $n$ draws. The urn scheme is run for a random number of draws $D_{t}$ to obtain a configuration $U_{D_{t}}$, where $\mathbb{P}(D_{t}=\ell)=d_{\ell}^{\theta}(t),~{}\ell\geq 1,$ and $D_{t}$ is as in (16) and $d_{\ell}^{\theta}(t)$ as in (18). The urn is then split and two urns are run independently for $n-D_{t}$ additional draws, for $n>D_{t}$, both beginning at $U_{D_{t}}$. This produces two random partitions $\eta$ and $\tilde{\eta}$, induced by the two resulting urn configurations. Conditional on the partition $\omega$ obtained at step $D_{t}$, these partitions are independent and with common distribution (24), namely
@@
$\mathbb{P}(U_{n}=\eta,U^{\prime}_{n}=\eta^{\prime})=\sum_{\ell=1}^{\infty}d_{\ell}^{\theta}(t)\sum_{|\omega|=\ell}\mathbb{E}_{\alpha,\theta}\big{[}\vec{P}_{\omega}\big{]}\mathbb{P}(\eta|\omega)\mathbb{P}(\eta^{\prime}|\omega),$
@@
$\sum_{\ell=1}^{\infty}d_{\ell}^{\theta}(t)p_{\ell}(x,y)\mathrm{PD}_{\alpha,\theta}(dx)\mathrm{PD}_{\alpha,\theta}(dy).$
This probabilistic construction shows that the conditional distribution of $X(t)$ given $X(0)$, where $X$ is the two-parameter diffusion, is the same as (26), at a fixed time $t$.
@@ Appendix A Appendix
by simply applying $L_{\alpha,\theta}$. If $a_{1}(\eta)>0$ $P_{\eta}$ can be recursively expressed as linear combinations of functions in ${\cal C}$ which can then be acted on by the differential operator (1). Recall that $P_{\eta}$ is exchangeable in the elements of $\eta$ so it is possible to rearrange the elements so that the $a_{1}(\eta)$ singletons are the last entries of $\eta$. Suppose that $\eta_{d}=1$. Then we use the notation that $\eta^{-}=\eta-e_{d}=(\eta_{1},\ldots,\eta_{d-1})$, where the final component of $\eta$ is removed. Consider (9) with $i=d$, then
@@
MR acknowledges support of MUR - Prin 2022 - Grant no. 2022CLTYP4, funded by the European Union - Next Generation EU. YZ was supported by grants NSFC117015170 and RDF-22-01-013.